Suppose A and B are events where P(A)=0.4, P(B)=0.5, and P(A and B)=0.1. So,
P(B|A)=?
P(A|B)=?
P(A or B)=? (if A and B are independent)
And why?

Answer by stanbon(75887)

Suppose A and B are events where P(A)=0.4, P(B)=0.5, and P(A and B)=0.1. So,
P(B|A)= P(A and B)/P(A) = 0.1/0.4 = 1/4
---------------------------------------------
P(A|B)= P(A and B)/P(B) = 0.1/0.5 = 1/5
---------------------------------------------
P(A or B)= P(A)*P(B) (if A and B are independent)
-----------------------
Cheers,
Stan H.